Find the equation in the from ax^2+bx+c=0 whose sum and product roots are respectively 3/2, —5/2.


1
Expert's answer
2021-03-09T08:08:04-0500

By Vieta's formula for a quadratic polynomial we have


"x_1+x_2=-\dfrac{b}{a}"

"x_1\cdot x_2=\dfrac{c}{a}"

Given


"x_1+x_2=\dfrac{3}{2}=-\dfrac{b}{a}"


"x_1\cdot x_2=-\dfrac{5}{2}=\dfrac{c}{a}"

Let "a=2." Then "b=-3, c=-5."

The equation is


"2x^2-3x-5=0"